# Generalizable Articulated Object Reconstruction from Casually Captured RGBD Videos

### v1

@@ C.2 Variance of joint estimation results
Joint | Methods | Axis | Position | Type | State | Failure |
Revolute | Ours | 0.300.03 | 0.130.01 | 0.150.02 | 0.250.00 | 0.060.00 |
Ours w/o Refine | 1.150.08 | 0.260.01 | 0.750.06 | 0.630.04 | 0.040.00 | |
Ours w/o Coarse | 1.260.01 | 0.340.00 | 0.790.03 | 0.690.01 | 0.020.00 | |
Ours w/o Segment | 0.380.03 | 0.170.00 | 0.190.02 | 0.330.00 | 0.060.00 | |
Prismatic | Ours | 0.230.03 | N.A. | 0.060.03 | 0.050.01 | 0.000.00 |
Ours w/o Refine | 0.290.06 | N.A. | 0.110.04 | 0.100.02 | 0.000.00 | |
Ours w/o Coarse | 0.480.01 | N.A. | 0.080.02 | 0.120.00 | 0.030.00 | |
Ours w/o Segment | 0.250.03 | N.A. | 0.080.02 | 0.070.00 | 0.000.00 |
To further evaluate the stability of our method, we conduct experiments on the S-Dataset with 10 different seeds. The mean and standard deviation of the final results are shown in Tab. C.4. We can find that our method consistently outperforms the ablated version of our method. Besides, by comparing our results and ours w/o Refinement results, we can find that the refinement module can reduce the variance of the prediction.
@@ C.3 Moving map segmentation results
Ours | 0.59653 |
Ours w/o Refine (MonST3R) | 0.61632 |
Ours w/o Coarse | 0.42545 |
Ours w/o Segment | 0.40310 |
Moving map segmentation is an important intermediate results of our method. We evaluate moving map segmentation using mIOU metric. From the quantitative results in Tab. C.5, we find our method achieves comparable results to MonST3R [55], though our pipeline ignores the inner parts of the object, which could be moving part in the video. Without coarse prediction as initialization, our method is hard to optimize for very accurate moving map segmentation. Similarly, without automatic mask generation, our pipeline cannot optimize moving map of the video efficiently, resulting in worse moving map segmentation of the video.
@@ C.4 Camera pose estimation results
Methods | Rotation Error | Position Error |
Ours | 0.08066 | 0.09246 |
@@
Ours w/o Refine | 0.11814 | 0.10918 |
Ours w/o Coarse | 0.11073 | 0.16544 |
Ours w/o Segment | 0.11547 | 0.16892 |
There are already several dynamic scene reconstruction works that can infer camera poses of a video sequence. This comparison aims to demonstrate why not directly use camera estimation results from existing dynamic scene reconstruction models. We compare our camera pose prediction results with two very recent works MonST3R [55] and CUT3R [42]. We transform all the camera poses to the camera coordinate of the first video frame. We compare both camera rotation error and translation error on our dataset. From the results shown in Tab. C.6, we can find that our method can produce more accurate camera pose estimation, particularly camera orientation, against recent works on dynamic scene understanding.
@@ C.5 Moving Map Prediction Choice
In our coarse prediction module, we use MonST3R [55] to generate a moving map of the input video. A natural question is, why not use an off-the-shelf motion segmentation method? Here we compare the output of moving map prediction results from MonST3R and Segment Any Motion [7], which is a very recent work on motion segmentation on videos. In Fig. C.2 we can find that Segment Any Motion produces much larger error prediction results compared to MonST3R. Segment Any Motion first computes point tracks of the video, and selects moving points based on point tracks. Then, it uses these points to prompt Segment Anything 2 to generate the moving map of the video. But in this case, the moving points lie around the edge of the door of this dishwasher. Thus, some points prompt selecting the dishwasher’s top as the moving part instead of the front door. This small position error will lead to a large error in moving map segmentation. In contrast, MonST3R computes the moving map based on optical flow and estimated camera poses. It subtracts the optical flow caused by camera motion from the final optical flow to determine which part of the video is moving. This strategy turns out to be much robust to point-track based motion segmentation methods.
C.6 Video segmentation strategy
Automatic part segmentation on the video plays a vital role in our pipeline. We find that automatic part segmentation performs better on the reversed video. Most videos in our dataset are about opening the door or drawer. In this case, some movable parts are difficult to segment out due to the very high similarity of texture of every part of the object. But when the door is opened or the drawer is pulled out, it’s much easier to distinguish the movable part from the other parts of the object. Fig. C.3 illustrates the segmentation results on a normal and reversed video.
@@ C.7 Additional real-world results
In the main paper we demonstrate the results of reconstructing real articulated objects with revolute joint. Here we demonstration reconstructing articulated objects with prismatic joints in Fig. C.4. Articulate Anything still struggles with retrieving correct object in Storage case, while RSRD fails to crop and cluster the object from the environment for the Drawer case.
@@ C.8 Failure case analysis
Our pipeline relies heavily on automatic part segmentation, since the refinement module is built upon part representation. We summarize three common failure modes in Fig. C.5. The first row shows that auto part segmentation may not identify newly observed parts in the video. It combines the interior of the drawer with its front face. When computing the Chamfer distance for optimization, the interior parts cannot find correspondence on the object surface points cloud, resulting in inaccurate estimation. The second row shows that auto part segmentation may not segment parts accurately. In this case, a part of the storage body is segmented out with the door. This part is static while the door is moving. This will also introduce noise to the Chamfer distance. The third row illustrates that auto part segmentation may not segment parts consistently. The lid of the trash can belongs to different segments in the video, shown in two different colors. Our pipeline will filter our parts that do not appear in the first frame. In this case, the trash can lid will be considered as newly observed parts and filtered out in the later sequence of the video, losing important information for joint estimation.

### v2

@@ C.2 Variance of Joint Estimation Results
| Joint | Methods | Axis(rad) | Position(m) | Type | State(rad or m) | Failure |
| Revolute | iTACO (Ours) | 0.300.03 | 0.130.01 | 0.150.02 | 0.250.00 | 0.060.00 |
| iTACO w/o Refine | 1.150.08 | 0.260.01 | 0.750.06 | 0.630.04 | 0.040.00 | |
| iTACO w/o Coarse | 1.260.01 | 0.340.00 | 0.790.03 | 0.690.01 | 0.020.00 | |
| iTACO w/o Segment | 0.380.03 | 0.170.00 | 0.190.02 | 0.330.00 | 0.060.00 | |
| Prismatic | iTACO | 0.230.03 | N.A. | 0.060.03 | 0.050.01 | 0.000.00 |
| iTACO w/o Refine | 0.290.06 | N.A. | 0.110.04 | 0.100.02 | 0.000.00 | |
| iTACO w/o Coarse | 0.480.01 | N.A. | 0.080.02 | 0.120.00 | 0.030.00 | |
| iTACO w/o Segment | 0.250.03 | N.A. | 0.080.02 | 0.070.00 | 0.000.00 |
To further evaluate the stability of our method, we conduct experiments on the S-Dataset with 10 different seeds. The mean and standard deviation of the final results are shown in Tab.˜C.4. We can find that our method consistently outperforms the ablated version of our method. Besides, by comparing our results and ours w/o Refinement results, we can find that the refinement module can reduce the variance of the prediction.
@@ C.3 Moving Map Segmentation Results
| iTACO (Ours) | 0.59653 |
| iTACO w/o Refine (MonST3R) | 0.61632 |
| iTACO w/o Coarse | 0.42545 |
| iTACO w/o Segment | 0.40310 |
Moving map segmentation is an important intermediate results of our method. We evaluate moving map segmentation using mIOU metric. From the quantitative results in Tab.˜C.5, we find our method achieves comparable results to MonST3R [82], though our pipeline ignores the inner parts of the object, which could be moving part in the video. Without coarse prediction as initialization, our method is hard to optimize for very accurate moving map segmentation. Similarly, without automatic mask generation, our pipeline cannot optimize moving map of the video efficiently, resulting in worse moving map segmentation of the video.
@@ C.4 Camera Pose Estimation Results
| Methods | Rot. Error(rad) | Pos. Error(m) |
@@
| iTACO (Ours) | 0.08066 | 0.09246 |
| iTACO w/o Refine | 0.11814 | 0.10918 |
| iTACO w/o Coarse | 0.11073 | 0.16544 |
| iTACO w/o Segment | 0.11547 | 0.16892 |
There are already several dynamic scene reconstruction works that can infer camera poses of a video sequence. This comparison aims to demonstrate why not directly use camera estimation results from existing dynamic scene reconstruction models. We compare our camera pose prediction results with two very recent works MonST3R [82] and CUT3R [57]. We transform all the camera poses to the camera coordinate of the first video frame. We compare both camera rotation error and translation error on our dataset. From the results shown in Tab.˜C.6, we can find that our method can produce more accurate camera pose estimation, particularly camera orientation, against recent works on dynamic scene understanding. We further evaluate how camera pose estimation will affect the final joint parameter estimation results on S-Dataset. In Tab.˜C.7, we find that using camera poses estimated by MonST3R will lead to a large degradation in the performance.
| Joint | Methods | Axis(rad) | Position(m) | Type(%) | State(rad or m) | Failure(%) |
| Revolute | iTACO (Ours) | 0.320.56 | 0.130.25 | 15.90 | 0.250.46 | 6.81 |
| iTACO w/ MonST3R camera | 0.740.58 | 0.270.25 | 51.11 | 0.390.43 | 6.67 | |
| Prismatic | iTACO | 0.240.33 | N.A. | 10.34 | 0.080.22 | 0 |
| iTACO w/ MonST3R camera | 0.680.31 | N.A. | 3.33 | 0.150.10 | 0 |
@@ C.5 Moving Map Prediction Choice
In our coarse prediction module, we use MonST3R [82] to generate a moving map of the input video. A natural question is, why not use an off-the-shelf motion segmentation method? Here we compare the output of moving map prediction results from MonST3R and Segment Any Motion [10], which is a very recent work on motion segmentation on videos. In Fig.˜C.2 we can find that Segment Any Motion produces much larger error prediction results compared to MonST3R. Segment Any Motion first computes point tracks of the video, and selects moving points based on point tracks. Then, it uses these points to prompt Segment Anything 2 [39] to generate the moving map of the video [10]. But in this case, the moving points lie around the edge of the door of this dishwasher. Thus, some points prompt selecting the dishwasher’s top as the moving part instead of the front door. This small position error will lead to a large error in moving map segmentation. In contrast, MonST3R computes the moving map based on optical flow and estimated camera poses. It subtracts the optical flow caused by camera motion from the final optical flow to determine which part of the video is moving. This strategy turns out to be much robust to point-track based motion segmentation methods.
C.6 Discussion on Automatic Video Segmentation Strategy
Automatic part segmentation on the video plays a vital role in our pipeline. We find that automatic part segmentation performs better on the reversed video. Most videos in our dataset are about opening the door or drawer. In this case, some movable parts are difficult to segment out due to the very high similarity of texture of every part of the object. But when the door is opened or the drawer is pulled out, it’s much easier to distinguish the movable part from the other parts of the object. Fig.˜C.3 illustrates the segmentation results on a normal and reversed video.
Besides, we test our method using ground truth mesh-level segmentation on the input video on S-Dataset. An image is rendered from an object comprising several meshes. The mesh-level segmentation identifies different regions that are rendered from different meshes. The results in Tab.˜C.8 show that at least using automatic part segmentation is comparable to using ground truth mesh-level segmentation.
| Joint | Methods | Axis(rad) | Position(m) | Type(%) | State(rad or m) | Failure(%) |
| Revolute | iTACO (Ours) | 0.320.56 | 0.130.25 | 15.90 | 0.250.46 | 6.81 |
| iTACO w/ mesh-level seg | 0.340.58 | 0.150.27 | 17.77 | 0.200.40 | 6.67 | |
| Prismatic | iTACO | 0.240.33 | N.A. | 10.34 | 0.080.22 | 0 |
| iTACO w/ mesh-level seg | 0.210.36 | N.A. | 10.00 | 0.030.13 | 0 |
@@ C.7 Additional Real-World Results
In the main paper, we demonstrate the results of building interactable digital twins of real articulated objects with revolute joints. Here we demonstrate reconstructing articulated objects with prismatic joints in Fig.˜C.4. Articulate Anything still struggles with retrieving the correct object in the Storage case, while RSRD fails to crop and cluster the object from the environment for the Drawer case.
@@ C.8 Failure Case Analysis
Our pipeline relies heavily on automatic part segmentation, since the refinement module is built upon part representation. We summarize three common failure modes in Fig.˜C.5. The first row shows that auto part segmentation may not identify newly observed parts in the video. It combines the interior of the drawer with its front face. When computing the Chamfer distance for optimization, the interior parts cannot find correspondence on the object surface points cloud, resulting in inaccurate estimation. The second row shows that auto part segmentation may not segment parts accurately. In this case, a part of the storage body is segmented out with the door. This part is static while the door is moving. This will also introduce noise to the Chamfer distance. The third row illustrates that auto part segmentation may not segment parts consistently. The lid of the trash can belongs to different segments in the video, shown in two different colors. Our pipeline will filter out parts that do not appear in the first frame. In this case, the trash can lid will be considered as newly observed parts and filtered out in the later sequence of the video, losing important information for joint estimation.